OpenStudy (anonymous):

Can someone help me? In isosceles triangle MNO, MN=6x+40, NO=4x+104, and OM=7x-23. If MN=OM, what is the perimeter of triangle MNO, in units?

OpenStudy (anonymous):

If MN=OM then you set up the problem like 6x+40=7x-23. 63=x. Then just plug 63 in for x in each equation. 6(63)+40=418, 4(63)+104=356, and 7(63)-23=418. Add them together to find the perimeter and you get 1192 units.